OpenStudy (anonymous):

how do you write y+3=4(x-1) in standard form

OpenStudy (anonymous):

Let's start by simplifying this equation as much as possible.

OpenStudy (anonymous):

Let's first distribute on the right side the 4(x-1) to be 4x-4

OpenStudy (anonymous):

Now we have, y + 3 = 4x + 4 Let's now combine the constant terms by either subtracting 4 on both sides, or subtracting 3 on both sides. Let's subtract 3 from both sides so we don't deal with negatives, which for making mistakes, we should avoid as much as possible. y + 3 - 3= 4x + 4 - 3 -> y = 4x + 1

OpenStudy (anonymous):

Now that we have y = 4x + 1, let's see what Standard Form looks like and what we need to do to rearrange this right. Standard Form is Ax + By + C = 0 Ok so let's set it like this: 4x + 1 = y then subtract y from both sides 4x + 1 - y = 0 switch the -y and 1 around 4x - y + 1 = 0 Ok there we go!! Now we know... A = 4 B = -1 C = 1 Our answer is 4x - y + 1 = 0

OpenStudy (anonymous):

thank you!